# NativQA Framework: Enabling LLMs with Native, Local, and Everyday Knowledge

### v1

@@ 5 Related Work
5.1 Local and Cultural Knowledge in LLMs
@@
LLMs have consistently demonstrated strong performance across a wide range of NLP tasks, with numerous efforts evaluating their capabilities on standard benchmarks Achiam et al. (2023); Touvron et al. (2023); Bubeck et al. (2023); Bang et al. (2023); Ahuja et al. (2023); Hendy et al. (2023); Abdelali et al. (2024). However, despite their scale and generalization abilities, LLMs often underperform on tasks that require culturally or locally grounded knowledge, particularly in low-resource languages and dialectal varieties Pawar et al. (2024). This limitation has sparked critical questions around the need for more culturally inclusive models. Recent work emphasizes the importance of enhancing LLMs to understand, reason about, and generate content grounded in everyday experiences, linguistic forms, and societal norms of diverse user communities Li et al. (2024b, a); Shi et al. (2024). A culturally inclusive model should be capable of interpreting how topics such as healthcare, education, or cuisine are framed in local contexts across different languages and dialects. Yet, current models often fall short in medium- to low-resource settings due to limited exposure to region-specific expressions, indigenous knowledge, and local conversations  Myung et al. (2024); Chiu et al. (2024).
5.2 Cultural, Information-Seeking, and Multilingual Resources
@@
Focusing on cultural aspects, there have been efforts to develop datasets that benchmark and fine-tune LLMs for diverse languages and regions Arora et al. (2024); Shi et al. (2024); Li et al. (2024b). In addition to cultural knowledge, there have also been initiatives to build resources that reflect everyday information-seeking behavior Myung et al. (2024). Efforts targeting everyday information-seeking queries include the development of mono- and multilingual resources by collecting data from diverse sources such as Wikipedia Yang et al. (2018); Kwiatkowski et al. (2019), Google Search QA Khashabi et al. (2021), Reddit forums Fan et al. (2019), and question–answer pairs written directly by native speakers Clark et al. (2020). Other approaches include combining native and machine-translated content Arora et al. (2024); Longpre et al. (2021); Asai et al. (2021), as well as generating QA data using LLMs Putri et al. (2024); Li et al. (2024b).
5.3 Motivation for NativQA framework
@@
The above studies make significant contributions toward incorporating native language questions and local cultural content. Early efforts such as TyDi QA Clark et al. (2020), BLEnD Myung et al. (2024), and CaLMQA Arora et al. (2024) expanded the coverage of culturally specific knowledge. Search engine-based datasets (e.g., NQ Kwiatkowski et al. (2019), MS MARCO Nguyen et al. (2016), and GooQA Khashabi et al. (2021)) demonstrated scalability and real-world relevance, although they often focus on a single dominant language. Despite this progress, the field still lacks a unified framework that jointly addresses multilinguality, location specificity, and scalability. NativQA framework sets itself apart by combining scalability with cultural specificity: it leverages real-world search queries and results to collect data at scale, while explicitly capturing native phrasing and culturally grounded contexts in underrepresented languages. This blend of scale and cultural relevance makes the NativQA framework a uniquely valuable tool for data collection that supports the training and evaluation of large language models (LLMs). As LLM development progresses, frameworks like NativQA fill a critical gap by providing native QA data essential for truly localized AI – an aspect only partially addressed by prior datasets.
@@ 6 Conclusions and Future Work
In this study, we propose the NativQA framework, which a scalable, language-independent framework to build culturally and regionally aligned QA datasets in native languages. It utilized search engines capabilities to collect questions and answers based on seed queries that native users ask. Based on the seed queries and target search locale one can collect a large number of QA pairs relatively quickly. The framework has already been used in published studies. We additionally evaluated its utility for another 30 locations from different countries and languages. Its functionality for multi-modalities can enable further research in this area and caching mechanism will help in reducing the search engines API cost. In future we aim to extend NativQA framework for other modalities such as audio/video and extensively evaluate for more under-representative locations.
@@ 7 Limitations
The proposed NativQA framework enables the development of datasets with cultural and native information for everyday knowledge. The NativQA framework currently uses human-in-the-loop processes in different phases such as seed query creation and manual revision of QA pairs. Though it also supports LLMs for seed query generation and QA validation, however, it is important to have a human-in-the-loop setup of the LLMs outputs. The framework does not include a built-in annotation platform for manual QA verification. However, any annotation platform (e.g., Pybossa [^2] https://github.com/Scifabric/pybossa) can be adapted for this purpose.
@@ Appendix C Distribution of Additional Dataset
In Figure A2, we show the distribution of the dataset across major cities and languages. The QA dataset covers seven major languages—Arabic, English, Hindi, Bangla, Turkish, Assamese, and Nepali—capturing a rich linguistic landscape spanning 39 global cities. Arabic QA pairs come from a variety of MENA countries, including Egypt, Iraq, Lebanon, and Qatar. English content is drawn from both Western countries (e.g., the USA and Canada) and multilingual regions such as Qatar and Bangladesh. South Asian languages like Hindi, Bangla, Assamese, and Nepali are represented through key urban centers including Delhi, Dhaka, Assam, Kolkata, and Kathmandu.
@@
Appendix D Annotation Guideline for Selecting the Best Answer
D.1 Objective
The purpose of this task is to evaluate two possible answers for a given question and select the one that is more accurate, clear, and well-structured. Each question will have two answer options, and your task is to choose the better answer based on the evaluation criteria provided below.
D.2 Annotation Task
For each question, you will see two answer options. Carefully compare them and select the one that best meets the following criteria:
@@ D.2 Annotation Task
    Accuracy The answer must be factually correct and free from misinformation. It should properly address the question and not introduce unrelated details.
@@
    Clarity & Readability The answer should be grammatically correct, well-structured, and easy to read. It should avoid ambiguity, repetition, or fragmented sentences. The information should be presented in a logical and coherent way.
@@
    Completeness & Relevance The answer should provide enough relevant information to fully respond to the question. It should avoid unnecessary details that do not contribute to answering the question. If an answer is too short or vague, it may be less useful than a more complete response.
@@
D.3 Annotation Instructions
  • Carefully read the question and both answer options.
  • Evaluate both answers based on the three main criteria:
@@ D.3 Annotation Instructions
    • Completeness & Relevance (Does it fully answer the question?)
  • Choose the better answer based on these criteria.
@@
  • If both answers are of equal quality, select the one that is more precise and easier to understand.
@@
  • If neither answer is acceptable (e.g., both contain major errors), select “neither” and leave a comment explaining why.
D.4 Special Cases & Edge Cases
[…]
    If both answers contain incorrect information → Select “Neither” and provide feedback.
[…]
    If one answer is correct but unclear, and the other is incorrect but well-written → Prefer the correct answer, and edit.
[…]
@@ D.4 Special Cases & Edge Cases
    If both answers are equally valid → Select the one that is clearer and more structured.
[…]
    If neither answer makes sense → Select “Neither” and provide feedback.

### v2

@@ 5 Related Work
Local and Cultural Knowledge in LLMs.
@@
LLMs excel on standard NLP benchmarks Achiam et al. (2023); Touvron et al. (2023); Bubeck et al. (2023); Bang et al. (2023); Ahuja et al. (2023); Hendy et al. (2023); Abdelali et al. (2024), yet often struggle with tasks requiring culturally grounded knowledge, especially in low-resource languages and dialects Pawar et al. (2024). This gap highlights the need for more inclusive models. Recent studies demonstrate the importance of enabling LLMs to understand and generate content reflective of diverse linguistic forms, social norms, and local experiences Li et al. (2024b, a); Shi et al. (2024). A culturally inclusive model should be capable of understanding topics like healthcare, education, or cuisine within local contexts. However, current LLMs frequently fall short in capturing region-specific expressions and indigenous knowledge Myung et al. (2024); Chiu et al. (2024).
Cultural, Information-Seeking, and Multilingual Resources.
@@
Focusing on cultural aspects, there have been efforts to develop datasets that benchmark and fine-tune LLMs for diverse languages and regions Arora et al. (2024); Shi et al. (2024); Li et al. (2024b). In addition to cultural knowledge, there have also been initiatives to build resources that reflect everyday information-seeking behavior Myung et al. (2024). Efforts targeting everyday information-seeking queries include the development of mono- and multilingual resources by collecting data from diverse sources such as Wikipedia Yang et al. (2018); Kwiatkowski et al. (2019), Google Search QA Khashabi et al. (2021), Reddit forums Fan et al. (2019), and question–answer pairs written directly by native speakers Clark et al. (2020). Other approaches include combining native and machine-translated content Arora et al. (2024), and generating QA using LLMs Putri et al. (2024); Li et al. (2024b).
Motivation for NativQA framework.
@@
Prior work has advanced the inclusion of native language questions and culturally specific content. Datasets like TyDi QA Clark et al. (2020), BLEnD Myung et al. (2024), and CaLMQA Arora et al. (2024) broaden cultural coverage, while search engine-based datasets such as NQ Kwiatkowski et al. (2019), MS MARCO Nguyen et al. (2016), and GooQA Khashabi et al. (2021) offer scalability and real-world relevance, albeit primarily in dominant languages. Despite these advances, a unified framework addressing multilinguality, location specificity, and scalability remains missing. NativQA framework bridges this gap by enabling large-scale data collection from real-world search queries with culturally grounded, native phrasing in underrepresented languages.
@@ 6 Conclusions and Future Work
In this study, we propose the NativQA framework, which a scalable, language-independent framework to build culturally and regionally aligned QA datasets in native languages. It utilizes search engines capabilities to collect QA based on seed queries that native users ask. Based on the seed queries and target search locale one can collect a large number of QA pairs relatively quickly. The framework has already been used in published studies. We additionally evaluated its utility for another 30 locations from different countries and languages. Its functionality for multi-modalities can enable further research in this area and caching mechanism will help in reducing the search engines API cost. In future we aim to extend NativQA framework for other modalities such as audio/video and evaluate for more under-representative locations.
@@ 7 Limitations
The proposed NativQA framework enables the development of datasets with cultural and native information for everyday knowledge. The NativQA framework currently uses human-in-the-loop processes in different phases such as seed query creation and manual revision of QA pairs. Though it also supports LLMs for seed query generation and QA validation, however, it is important to have a human-in-the-loop setup to verify the LLMs outputs. The framework does not include a built-in annotation platform for manual QA verification. However, any annotation platform (e.g., Pybossa [^2] https://github.com/Scifabric/pybossa) can be adapted for this purpose.
@@ Appendix C Distribution of Additional Dataset
Language Location Train Dev Test Total
Arabic Abu Dhabi, UAE 4,896 692 1,406 6,994
Algiers, Algeria 15,198 2,149 4,364 21,711
Baghdad, Iraq 17,910 2,533 5,143 25,586
Beirut, Lebanon 17,748 2,510 5,096 25,354
Cairo, Egypt 8,631 1,221 2,478 12,330
Damascus, Syria 7,902 1,117 2,269 11,288
Doha, Qatar 3,649 492 988 5,129
Gaza, Palestinian 3,975 562 1,142 5,679
Khartoum, Sudan 3303 467 948 4,718
Kuwait City, Kuwait 21,676 3,066 6,224 30,966
Manama, Bahrain 17,599 2,489 5,053 25,141
Muscat, Oman 3,479 492 999 4,970
Nouakchott, Mauritania 13,520 1,912 3,882 19,314
Rabat, Morocco 17,341 2,453 4,979 24,773
Riyadh, Saudi Arabia 4,150 587 1,191 5,928
Sanaa, Yemen 3,373 477 968 4,818
Tripoli, Libya 3,300 467 947 4,714
Tunis, Tunisia 10,352 1,464 2,972 14,788
Assamese Assam, India 1,131 157 545 1,833
Bangla Dhaka, Bangladesh 7,018 953 1,521 9,492
Kolkata, India 6,891 930 2,146 9,967
English California, USA 1,065 151 306 1,522
Dhaka, Bangladesh 4,761 656 1,113 6,530
Doha, Qatar 8,212 1,164 2,322 11,698
Florida, USA 777 110 223 1,110
Georgia, USA 816 115 234 1,165
Hawaii, USA 788 112 226 1,126
Illinois, USA 877 124 252 1,253
Massachusetts, USA 842 119 242 1,203
Michigan, USA 796 113 228 1,137
New York, USA 4,518 623 1,313 6,454
North Carolina, USA 809 115 232 1,156
Ohio, USA 862 122 247 1,231
Ontario, Canada 743 105 213 1,061
Pennsylvania, USA 786 111 226 1,123
Quebec, Canada 720 102 207 1,029
Texas, USA 847 120 243 1,210
Washington, USA 874 124 251 1,249
Hindi Delhi, India 9,288 1,286 2,745 13,319
Nepali Kathmandu, Nepal 561 561
Turkish Istanbul, Turkey 3,527 483 1,218 5,228
Total 234,950 33,045 67,863 335,858
Table A1: Statistics of the data collected using NativQA Framework. – Only testing split due to limited dataset size.
In Figure A2, we show the distribution of the dataset across major cities and languages. The QA dataset covers seven languages—Arabic, English, Hindi, Bangla, Turkish, Assamese, and Nepali—capturing a rich linguistic landscape spanning 39 global cities. Arabic QA pairs come from a variety of MENA countries, including Egypt, Iraq, Lebanon, and Qatar. English content is drawn from both Western countries (e.g., the USA and Canada) and multilingual regions such as Qatar and Bangladesh. South Asian languages like Hindi, Bangla, Assamese, and Nepali are represented through key urban centers including Delhi, Dhaka, Assam, Kolkata, and Kathmandu. Moreover, we present the details of the collected data using NativQA framework in Table A1. We split the data into training (70%), development (10%), and test (20%) sets.
@@
Appendix D Annotation Guideline
D.1 Manual QA validation - Selecting Best Answer
The purpose of this task is to evaluate two possible answers for a given question and select the one that is more accurate, clear, and well-structured. Each question will have two answer options, and the task is to choose the better answer based on the evaluation criteria discussed below.
D.1.1 Annotation Task
For each question, there are two answer options. The task is to carefully compare them and select the one that best meets the below criteria:
@@ D.1.1 Annotation Task
    Accuracy. The answer must be factually correct and free from misinformation. It should properly address the question and not introduce unrelated details.
@@
    Clarity & Readability. The answer should be grammatically correct, well-structured, and easy to read. It should avoid ambiguity, repetition, or fragmented sentences. The information should be presented in a logical and coherent way.
@@
    Completeness & Relevance. The answer should provide enough relevant information to fully respond to the question. It should avoid unnecessary details that do not contribute to answering the question. If an answer is too short or vague, it may be less useful than a more complete response.
@@
D.1.2 Annotation Instructions
  • The question and both answer options should be read carefully.
  • Both answers should be evaluated based on the following three main criteria:
@@ D.1.2 Annotation Instructions
    • Completeness & Relevance (Is the question fully and appropriately answered?)
  • The better answer should be selected based on these criteria.
@@
  • If the answers are of equal quality, the one that is more precise and easier to understand should be chosen.
@@
  • If neither answer is acceptable (e.g., both contain major errors), “neither” should be selected, and a comment explaining the reasoning should be provided.
D.1.3 Special Cases and Edge Cases
  • If one answer is factually correct but lacks clarity, while the other is well-written but factually incorrect, the correct answer should be preferred and revised for clarity if necessary.
@@ D.1.3 Special Cases and Edge Cases
  • If both answers are equally valid in terms of content, the answer that is clearer and better structured should be selected.
D.2 Manual QA validation - Answer Scoring.
The goal of this task is to assess and, if necessary, revise LLM-edited answers using the information provided in the associated source URLs. Each question includes an LLM-generated answer and its corresponding source. Annotators are expected to verify the quality of the answer based on the source content, and evaluate it according to the criteria outlined below.
Figure A3: An example of the annotation interface for QA annotation for LLM-based QA annotation.
D.2.1 Annotation Task
To assess the quality of LLM-edited answers, we used four different metrics described below. For each metric, assign a score from 1 (very poor) to 5 (excellent), following the descriptions below. Informativeness. Measures the extent to which the assessed text (answer) provides relevant, meaningful, complete (covers all important points), correct (factually accurate), and useful information in response to the given prompt or context. Highly informative content (answer) directly addresses key points; low-informative content (answer) is vague, incomplete, incorrect, or misses critical details.
  • 1 = Not informative: Lacks relevant or meaningful content; does not address the question/context; contains major errors or omissions.
  • 2 = Slightly informative: Provides minimal information; many important details are missing, unclear, or inaccurate.
  • 3 = Moderately informative: Some relevant details are present but lack depth, accuracy, or completeness.
  • 4 = Informative: Well-detailed; addresses most key points clearly, correctly, and thoroughly.
  • 5 = Very informative: Exceptionally detailed, insightful, accurate, and complete; fully addresses all aspects of the prompt/context.
Clarity. Assesses how clearly the assessed text (answer) conveys its meaning. Clear text is well-structured, concise (free from unnecessary information), and easy to understand, free from ambiguity, awkward phrasing, or grammatical errors.
  • 1 = Very unclear: Difficult or impossible to understand; contains major grammatical or structural problems; excessively wordy or redundant.
  • 2 = Somewhat unclear: Some meaning is discernible, but significant ambiguity, awkwardness, or unnecessary verbosity remains.
  • 3 = Neutral: Generally understandable but may require effort or contain minor issues, including some unnecessary length.
  • 4 = Clear: Easy to read and understand with minimal ambiguity; well-structured; mostly concise.
  • 5 = Very clear: Highly readable, precise, concise, and effortlessly understandable.
Plausibility. Refers to whether the answer makes sense and is reasonable in the given context. A plausible response is coherent, logical, correct (factually accurate), and consistent with the input (e.g., question, original text, source image/audio).
  • 1 = Not plausible at all: Response is illogical, unrelated, factually incorrect, or clearly incorrect.
  • 2 = Weakly plausible: Some connection, but contains major logical flaws, inaccuracies, or inconsistencies.
  • 3 = Moderately plausible: Partially logical, but some elements are incomplete, questionable, or partially incorrect.
  • 4 = Plausible: Reasonable, makes sense, factually correct, and is mostly consistent with the context.
  • 5 = Highly plausible: Fully logical, coherent, accurate, and strongly aligned with the context.
Faithfulness. Measures how accurately the assessed text reflects or stays true to the source information or reasoning. Faithful responses avoid adding unrelated details, omitting crucial information (completeness), or misrepresenting facts (correctness and honesty).
  • 1 = Not faithful at all: Largely or entirely unrelated to the source/context; misrepresents information; misses key points or adds incorrect information.
  • 2 = Weakly faithful: Some relevant aspects, but much is misleading, incomplete, or missing.
  • 3 = Moderately faithful: Captures some correct elements, but contains errors, omissions, or partial misrepresentation.
  • 4 = Faithful: Accurately represents the main points and intent of the source/context, without major errors or omissions.
  • 5 = Highly faithful: Fully preserves the original meaning/reasoning, covers all important aspects, and is free from errors or misleading information.
D.3 Annotation Instructions
  • Read carefully: Always review both the context and the text to be assessed as instructed in the platform.
  • Apply each metric independently: Judge informativeness, clarity, plausibility, and faithfulness separately for each item.
  • Be objective: Follow the criteria closely; avoid personal biases or preferences.
  • Consider the context: Assess whether the text meets the needs of the specific task (e.g., Does the answer address the question?).
  • Use full scale: Do not hesitate to assign low or high scores where appropriate. A score of 1 means very poor; 5 means excellent.
D.4 Annotation Interface
Figure A3 present the annotation interface used by annotators to evaluate the edited answers. The source article, shown on the right side of the interface, is used to verify the accuracy and completeness of the answer.